# DESI results and Dark Energy from QCD topological sectors

### v1

@@ Appendix B Few technical comments on the nature of the QCD-induced DE
1. The main claim of [8] is the presence of a linear correction, as expressed in Eq.(1). This constitutes a central element of the entire proposal, as one would naively expect a quadratic correction at small curvature, i.e., $\propto\kappa^{2}$. Indeed, the only difference between the two geometries $\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}$ and $\mathbb{R}^{3}\times\mathbb{S}^{1}$—assuming identical sizes for $\mathbb{S}^{1}_{\kappa^{-1}}$ and $\mathbb{S}^{1}$—is the small curvature $\sim\kappa^{2}$ of the hyperbolic space $\mathbb{H}^{3}_{\kappa}$, in contrast to the zero curvature of $\mathbb{R}^{3}$. Conventional locality arguments would therefore unambiguously suggest that any correction to the vacuum energy difference in Eq.(1) should be expressible solely in terms of the curvature, and thus involve only even powers of $\kappa$, i.e., $\propto\kappa^{2n}$. However, explicit calculations in [8] reveal a linear correction, contradicting this expectation.
[…]
@@
Therefore, the standard arguments based on locality are strongly violated by such configurations. As shown in the computations of [8], the linear correction $\sim\kappa$ is explicitly proportional to the holonomy defined in Eq. (30), a gauge-invariant observable that cannot be reduced to the local curvature. In other words, this correction arises from non-local configurations and cannot be expressed in terms of the local curvature $\sim\kappa^{2}$.
@@
3. The linear correction $\sim\kappa$ in $\Delta E_{\rm vac}$ can be traced, at a technical level, to the differing behavior of monopole configurations at large distances in the two backgrounds. In hyperbolic space $\mathbb{H}^{3}_{\kappa}$, monopole fields experience an exponential suppression at the characteristic scale $\kappa^{-1}$, whereas no such cutoff is present in flat space $\mathbb{R}^{3}$ [8].
@@
4. Another technical reason that makes these computations feasible is the conformal equivalence between $\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}$ and $\mathbb{R}^{4}$, which ensures that the subtraction procedure in Eq. (1) is well defined [8]. This is a crucial point because the number of zero modes (i.e., the moduli space) contributing to the path integral is identical in both geometries, and the expressions for their volumes coincide in the small $\kappa$ limit, given a fixed large radius. As a result, the computation yields an energy expression that is infrared finite and exhibits extensive scaling with volume, i.e., $E\propto\rho V$. This outcome is highly nontrivial in the context of non-perturbative calculations, where achieving extensivity typically requires summing over an infinite number of monopole configurations. For further details, see [8].
@@
6. This non-dispersive contact term cannot be removed by any UV renormalization procedures as it has the IR nature, see [9] for more details.

### v2

@@ Appendix B Few technical comments on the nature of the QCD-induced DE
1. The main claim of Zhitnitsky (2015) is the presence of a linear correction, as expressed in Eq.(1). This constitutes a central element of the entire proposal, as one would naively expect a quadratic correction at small curvature, i.e., $\propto\kappa^{2}$. Indeed, the only difference between the two geometries $\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}$ and $\mathbb{R}^{3}\times\mathbb{S}^{1}$—assuming identical sizes for $\mathbb{S}^{1}_{\kappa^{-1}}$ and $\mathbb{S}^{1}$—is the small curvature $\sim\kappa^{2}$ of the hyperbolic space $\mathbb{H}^{3}_{\kappa}$, in contrast to the zero curvature of $\mathbb{R}^{3}$. Conventional locality arguments would therefore unambiguously suggest that any correction to the vacuum energy difference in Eq.(1) should be expressible solely in terms of the curvature, and thus involve only even powers of $\kappa$, i.e., $\propto\kappa^{2n}$. However, explicit calculations in Zhitnitsky (2015) reveal a linear correction, contradicting this expectation.
2. This linear correction, $\sim\kappa$, which plays a central role in the present proposal, is generated by vacuum configurations with nontrivial holonomy. The holonomy is defined as $U(\mathbf{x})\equiv{\cal{P}}\exp\left(i\int_{0}^{{\cal{T}}}dx_{4}A_{4}(x_{4},\mathbf{x})\right)$, where $\mathcal{P}$ denotes path ordering. Specifically, we consider the Polyakov line evaluated at spatial infinity, i.e.
$L={\cal{P}}\exp\left(i\int_{0}^{{\cal{T}}}dx_{4}A_{4}(x_{4},|\mathbf{x}|\rightarrow\infty)\right).$ (B1)
The operator $\mathrm{Tr}L$ classifies the self-dual solutions that may contribute to the path integral at finite temperature, $T\equiv\mathcal{T}^{-1}$, including the low-temperature limit as $T\rightarrow 0$. In particular, for $SU(2)$ gauge group the holonomy
$\frac{1}{2}\mathrm{Tr}L=\cos(\pi\nu),$ (B2)
belongs to the group center $\frac{1}{2}\mathrm{Tr}L=\pm 1$ when $\nu$ assumes the integer values (trivial holonomy). The so-called confining value for the holonomy corresponds to $\nu=1/2$ when $\mathrm{Tr}L=0$ vanishes.
@@
Therefore, the standard arguments based on locality are strongly violated by such configurations. As shown in the computations of Zhitnitsky (2015), the linear correction $\sim\kappa$ is explicitly proportional to the holonomy defined in Eq. (B2), a gauge-invariant observable that cannot be reduced to the local curvature. In other words, this correction arises from non-local configurations and cannot be expressed in terms of the local curvature $\sim\kappa^{2}$.
@@
3. The linear correction $\sim\kappa$ in $\Delta E_{\rm vac}$ can be traced, at a technical level, to the differing behavior of monopole configurations at large distances in the two backgrounds. In hyperbolic space $\mathbb{H}^{3}_{\kappa}$, monopole fields experience an exponential suppression at the characteristic scale $\kappa^{-1}$, whereas no such cutoff is present in flat space $\mathbb{R}^{3}$ Zhitnitsky (2015).
@@
4. Another technical reason that makes these computations feasible is the conformal equivalence between $\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}$ and $\mathbb{R}^{4}$, which ensures that the subtraction procedure in Eq. (1) is well defined Zhitnitsky (2015). This is a crucial point because the number of zero modes (i.e., the moduli space) contributing to the path integral is identical in both geometries, and the expressions for their volumes coincide in the small $\kappa$ limit, given a fixed large radius. As a result, the computation yields an energy expression that is infrared finite and exhibits extensive scaling with volume, i.e., $E\propto\rho V$. This outcome is highly nontrivial in the context of non-perturbative calculations, where achieving extensivity typically requires summing over an infinite number of monopole configurations. For further details, see Zhitnitsky (2015).
@@
6. This non-dispersive contact term cannot be removed by any UV renormalization procedures as it has the IR nature, see Barvinsky and Zhitnitsky (2018) for more details.